Fermi National Laboratory


Accelerator Update for December 27, 2004 - January 3, 2005

Monday December 27
The day shift began with Operations monitoring store 3879, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

A Main Injector (MI) expert reported at 8:54 AM that all 18 RF stations were working.

At 1:30 PM, Linac technicians began replacing the power amplifier for a RF station (LRF5).

Due to an MI kicker problem, operators began making MI ready for an access at 12:31 PM. Operators also made Booster and two Meson areas ready for access.

LRF5 was up and running by 4:48 PM.

Operations terminated store 3879 at 5 PM and then immediately put the Tevatron (TeV) into shot setup.

Operations established store 3881 at 10:42 PM with an initial luminosity of 64.94E30. Pbar resumed stacking and Operations resumed sending beam to MiniBooNE and to MCenter.

Tuesday December 28
The midnight shift began with Operations monitoring store 3881, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations and a Recycler expert began setting up Recycler at 3:35 AM to take antiprotons from Pbar.

Pbar resumed stacking at 4:52 AM.

Recycler began stashing more antiprotons at 10:17 AM. They took three transfers and then stacking resumed at 10:29 AM.

At 1:41 PM, MiniBooNE suffered the first of many trips throughout the afternoon.

Around 11 PM, Pbar began having emittance problems that slowed stacking.

Wednesday December 29
The midnight shift began with Operations monitoring store 3881 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations terminated store 3881 at 6:03 AM.

Operations began preparing the TeV for an access at 7:30 AM to repair a kicker in TeV sector F17. Booster techs began some RF work.

Techs completed the Booster RF work at 10:08 AM. Beam to MiniBooNE and MCenter resumed at 10:18 AM and Pbar resumed stacking at 10:35 AM.

Operations put the TeV into shot setup at 1:05 PM.

At 3:09 PM, the TeV suffered a quench while loading antiprotons. The Pbar abort kicker caused the quench.

TeV experts began a study period at 9:27 PM.

Thursday December 30
The midnight shift began with Operations monitoring the Recycler stash, with Pbar stacking, with TeV conducting studies, and with MiniBooNE and MCenter taking beam.

Operations put the TeV into shot setup at 3:28 AM.

Operations established store 3887 at 5:49 AM with an initial luminosity of 64.79E30. Pbar resumed stacking and MiniBooNE and MCenter resumed taking beam.

At 9:26 AM, operators began preparing Booster for an access to replace RF station air filters.

The Booster access ended and stacking resumed at 10:58 AM.

At 3:02 PM, techs at CHL reported a power glitch that affected some equipment. MCenter and MTest also reported having a power problem. The Duty Electrician went to the Master Substation to investigate. A high voltage expert came in and reported that feeder 30 had faulted. He switched its breaker off. Since this feeder has no load, it will be repaired later.

Pbar's Debuncher water skid tripped off at 9:17 PM. LCW experts came in to help operators get everything running normally and Pbar resumed stacking at 10:50 PM.

Friday December 31
The midnight shift began with Operation monitoring store 3887 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations began sending beam to MTest at 12:30 AM.

Operations terminated store 3887 at 5:02 AM and put the TeV into shot setup twenty minutes later.

At 7:18 AM, Operations reported problems with a Linac RF station (KRF1). Operations eventually called in a tech.

At 7:44 AM, the TeV suffered a sector D1 quench during the squeeze. Pbar resumed stacking.

A Linac expert got KRF1 running and Pbar resumed stacking at 8:40 AM.

A TeV expert began a study period at 12:02 PM. A NuMI expert began a study a few minutes later.

Operations put the TeV into shot setup and established store 3892 at 6:33 PM. Its initial luminosity was 39.2E30.

At 10:06 PM, a Booster kicker (MKS03) tripped off and wouldn't reset. Since it was New Years Eve and the booster could still operate in most of its modes, the kicker was left off.

Saturday January 1
The New Year midnight shift began with Operations monitoring store 3892 and a Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

A Booster expert came in and had the kicker (MKS03) repaired and running by 9:31 AM.

Operations terminated store 3892 at noon and began shot setup at 12:14 PM.

Operations established store 3893 at 2:19 PM with an initial luminosity of 63.95E30. Pbar resumed stacking and Operations resumed sending beam to MiniBooNE and the Meson experiments.

The Recycler began stashing antiprotons at 5:56 PM. Stacking resumed at 6:48 PM.

At 9:32 PM, Operations reported that the village was without power. ComEd reported that they lost power lines due to a vehicle accident.

Sunday January 2
The midnight shift began with Operations monitoring store 3893 and a Recycler stash, with Pbar stacking, and with beam to MiniBooNE and MCenter.

Operations terminated store 3893 at 4:04 PM.

Operation put the TeV into shot setup at 5:13 PM.

A series of Booster problems, which began around 5:21 PM, put the shot setup on hold until 7:33 PM.

Operations established store 3896 with an initial luminosity of 80.0E30.

At 11:49 PM, Operations reported a power outage. Store 3896 aborted cleanly.

Monday January 3
The midnight shift began with Operations trying to recover from the power outage.

At 12:32 AM, a second power outage hit and the power remained off. High voltage experts restored site wide power at 2:10 AM by back feeding the site through the Kautz Rd. substation. Recovery began again.

ComEd reported at 2:56 AM that they had a relay switch failure. They had the relay replaced by 6 AM.

The Plan for Monday
The plan for today is to continue with machine checkouts and recovery.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/3/2005   email Fermilab